Question 6 1 pts Assume that you consume the same basket of goods (apples and shirts) each year given by the data below which includes the per-unit price of each item in 2015 and 2016. If 2015 is the base year, calculate the CPl in 2016. (please round your answer to the nearest one decimal place) Basket: 6 apples, 10 shirts apples shirts 2015 $2 2016 2.50 $%23 $20
Answer #1

Solution -

Consumer Price Index = Cost of Market Basket (Current) / Cost of Market Basket ( Base )

  

Cost Of Basket 2015 = 6 * 2 + 10 * 20

= 12 + 200

= 212

Cost Of Basket 2016 = 6 * 2.50+23*10

= 15 + 230

  = 245

CPI in 2016(using 2015 Base Year) = 245 / 212 * 100

= 1.1556*100

= 115.56

= 116

CPI in 2016(using 2015 Base Year) = 116